OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and has a slope of 2 in point-slope form.

OpenStudy (priyar):

do u know the formula?

OpenStudy (whpalmer4):

point-slope form: \[y-y_0 = m (x - x_0)\]where \(m\) is slope, and \((x_0,y_0)\) is the known point

OpenStudy (dayakar):

slope (m) =2 (x0 ,y0) =(1, 3) substitute these values in above equation

OpenStudy (marcelie):



OpenStudy (marcelie):

i thinks that the answer

OpenStudy (welshfella):

y - 3 = 2(x - 1) simplify

OpenStudy (whpalmer4):

The problem does ask for the answer in point-slope form...while \(y=2x+1\) is the equation of the solution, \(y - 3 = 2(x-1)\) is how I would answer when asked for "equation ... in point-slope form"